# Data Mixing Can Induce Phase Transitions in Knowledge Acquisition

### v1

@@ Appendix A Related Works
LLMs are typically trained on a vast amount of data that are rich in knowledge, and extensive studies have investigated how much knowledge LLMs can acquire from the training data. Pioneering studies [Petroni et al., 2019, Roberts et al., 2020, Da et al., 2021] demonstrate that LLMs can capture a substantial amount of knowledge, suggesting their potential as knowledge bases. To quantify the relationship between model size and knowledge storage, Allen-Zhu and Li [2024a] and Lu et al. [2024] discover a linear relationship between models’ knowledge capacity and their parameter count by training LLMs on data only containing fixed-format knowledge for sufficiently long horizons. Later, Nichani et al. [2025] formally proved this linear relationship. In contrast, this paper examines the data mixing scenario and demonstrates that this linear scaling can be disrupted when the knowledge-dense dataset is mixed with vast amounts of web-scraped data. Another important factor is the frequency of occurrence for knowledge.
@@
This paper identifies phase transitions in knowledge acquisition within data mixtures with respect to model size and mixing ratio. Some relevant observations can be found in previous papers, but we takes a more direct and systematic approach. Kandpal et al. [2023], Mallen et al. [2023], Sun et al. [2024] find that LLMs can perform poorly on low-frequency knowledge. Ghosal et al. [2024] show that frequency of knowledge in the pre-training data determines how well the model encodes the knowledge, which influences its extractability after QA fine-tuning. Taking a more microscopic view, Chang et al. [2024] insert a few pieces of new knowledge during training and track their loss. By fitting a forgetting curve, they conjecture that the model may fail to learn the knowledge if its frequency is lower than some threshold.
@@
Our findings also relate to prior observations on the memorization and forgetting behaviors of LLMs, but we explicitly characterize phase transitions in the context of data mixing. Carlini et al. [2023] show that memorization of training data follows a log-linear relationship with model size, the number of repetitions, and prompt length. Biderman et al. [2024] take a data point-level perspective and demonstrate that it is difficult to predict whether a given data point will be memorized using a smaller or partially trained model. By injecting a few new sequences into the training data, Huang et al. [2024] find that a sequence must be repeated a non-trivial number of times to be memorized. By examining training dynamics, Tirumala et al. [2022] observe that memorization can occur before overfitting and that larger models memorize faster while forgetting more slowly. From a theoretical perspective, Feldman [2020] prove that memorization of training labels is necessary to achieve near-optimal generalization error for long-tailed data distributions.
@@
LLM performance is significantly influenced by the mixing proportions of the training data from different domains. Our paper is related to a line of studies that optimize the mixing proportions by modeling LLM performance as a function of the mixing proportions [Liu et al., 2024, Kang et al., 2024, Ye et al., 2024, Ge et al., 2024]. However, their datasets can be highly heterogeneous even within a single domain (e.g., OpenWebText, Pile-CC) while we focus on mixing a uniform, knowledge-dense dataset into web-scraped data.
Appendix B Additional Experimental Results
B.1 Phase Transition in Mixing Ratio for Larger Models
@@
B.2 Additional Plots for Mitigation Strategies
@@
Figure 11: Additional plots for mitigation strategies.
B.3 Additional Results for Validating the Power-Law Relationship of Threshold Frequency and Model Size
In Figure 12, we relax the constraint of training on the same data mixture and investigate the overall trend between model size and $P_{\mathrm{thres}}$. We add the Llama-3 [Dubey et al., 2024] family, and evaluate both base and instruction-tuned models for all families, totaling 30 models. Interestingly, in Figure 12, log model size and $\log P_{\mathrm{thres}}$ also exhibit a linear relationship, with most models falling within the 95% confidence interval. We further use models from the OLMo [Groeneveld et al., 2024] family as a validation set, where predictions of the fitted power law closely match the ground truth.
Potential Application: Inferring the Size of Proprietary Models.
B.4 Detailed Performance on SynBio
@@
$N$ $\rho$ (%) $r$ Birth date Birth city University Major Employer Avg.
@@
$N$ $\rho$ (%) $r$ Training tokens (B) Birth date Birth city University Major Employer Avg.
@@
B.5 Detailed Downstream Performance
We employ the lm-eval-harness [Gao et al., 2024] codebase to evaluate the zero-shot performance on five downstream tasks, including LAMBADA [Paperno et al., 2016], ARC-E [Clark et al., 2018], PIQA [Bisk et al., 2020], SciQ [Welbl et al., 2017], and HellaSwag [Zellers et al., 2019], covering core capabilities such as text understanding, commonsense reasoning, and question answering. We compute the validation loss on about 50M tokens on a holdout set from the Pile or FineWeb-Edu. The detailed downstream performance and validation loss for applying the random subsampling strategy to SynBio and WikiBio are presented in Tables 2 and 3, respectively. Additionally, we report the detailed downstream results for applying CKM to WikiBio in Table 4.
@@ B.5 Detailed Downstream Performance
$N$ $\rho$ (%) $r$ LAMBADA ARC-E Sciq PIQA HellaSwag Avg. FineWeb-Edu val. loss
- - 00 $38.25_{(0.68)}$ $61.83_{(1.00)}$ $83.60_{(1.17)}$ $68.01_{(1.09)}$ $35.04_{(0.48)}$ $57.35$ $2.667$
$1.28$M $100$ $0.1$ $34.56_{(0.66)}$ $62.33_{(0.99)}$ $83.50_{(1.17)}$ $68.34_{(1.09)}$ $35.13_{(0.48)}$ $56.77(\downarrow 0.58)$ $2.668(\uparrow 0.001)$
$1.28$M $100$ $0.2$ $34.43_{(0.67)}$ $62.13_{(0.99)}$ $83.80_{(1.17)}$ $68.12_{(1.09)}$ $35.39_{(0.48)}$ $56.77(\downarrow 0.58)$ $2.668(\uparrow 0.001)$
$1.28$M $100$ $0.3$ $33.94_{(0.66)}$ $60.77_{(1.00)}$ $80.80_{(1.25)}$ $66.54_{(1.10)}$ $34.23_{(0.47)}$ $55.26(\downarrow 2.09)$ $2.722(\uparrow 0.054)$
$320$k $25$ $0.2$ $36.70_{(0.67)}$ $60.35_{(1.00)}$ $(82.70_{1.20})$ $67.74_{(1.09)}$ $34.76_{(0.48)}$ $56.45(\downarrow 0.90)$ $2.686(\uparrow 0.019)$
$640$k $50$ $0.2$ $36.58_{(0.67)}$ $60.61_{(1.00)}$ $83.30_{(1.18)}$ $66.65_{(1.10)}$ $34.53_{(0.47)}$ $56.33(\downarrow 1.02)$ $2.688(\uparrow 0.021)$
$720$k $56.25$ $0.2$ $35.61_{(0.67)}$ $60.94_{(1.00)}$ $83.00_{(1.19)}$ $67.14_{(1.10)}$ $34.54_{(0.47)}$ $56.25(\downarrow 1.10)$ $2.687(\uparrow 0.020)$
$800$k $62.5$ $0.2$ $35.20_{(0.67)}$ $60.48_{(1.00)}$ $83.40_{(1.20)}$ $66.54_{(1.10)}$ $34.45_{(0.47)}$ $56.01(\downarrow 1.34)$ $2.688(\uparrow 0.021)$
@@
$N$ $\rho$ (%) $r$ Training tokens (B) LAMBADA ARC-E Sciq PIQA HellaSwag Avg. Pile val. loss
@@
$2.56$M $100$ $0.2$ $64$ $53.68_{(0.69)}$ $51.47_{(1.03)}$ $81.00_{(1.24)}$ $68.77_{(1.08)}$ $35.91_{(0.48)}$ $58.17(\downarrow 2.01)$ $2.184(\uparrow 0.016)$
$2.56$M $100$ $0.4$ $24$ $52.38_{(0.70)}$ $51.47_{(1.03)}$ $80.70_{(1.25)}$ $68.17_{(1.09)}$ $34.95_{(0.48)}$ $57.53(\downarrow 2.65)$ $2.198(\uparrow 0.030)$
$1.28$M $50$ $0.2$ $64$ $54.71_{(0.69)}$ $52.86_{(1.02)}$ $81.30_{(1.23)}$ $68.99_{(1.08)}$ $35.48_{(0.48)}$ $58.67(\downarrow 1.51)$ $2.189(\uparrow 0.022)$
@@
$N$ $\rho$ (%) $r$ Training tokens (B) LAMBADA ARC-E Sciq PIQA HellaSwag Avg. Pile val. loss
@@
$277$k $100$ $0.1$ $32$ $50.77_{(0.70)}$ $48.95_{(1.03)}$ $80.80_{(1.25)}$ $66.43_{(1.10)}$ $33.16_{(0.47)}$ $56.02(\downarrow 1.56)$ $2.286(\uparrow 0.031)$
$277$k $100$ $0.15$ $32$ $49.12_{(0.70)}$ $49.66_{(1.03)}$ $81.80_{(1.22)}$ $66.38_{(1.10)}$ $32.84_{(0.47)}$ $55.96(\downarrow 1.62)$ $2.292(\uparrow 0.037)$
$277$k $100$ $0.2$ $20$ $49.37_{(0.70)}$ $49.87_{(1.03)}$ $79.70_{(1.27)}$ $65.40_{(1.11)}$ $33.08_{(0.47)}$ $55.48(\downarrow 2.10)$ $2.306(\uparrow 0.051)$
$69$k $25$ $0.1$ $32$ $48.63_{(0.70)}$ $50.59_{(1.03)}$ $81.00_{(1.24)}$ $66.49_{(1.10)}$ $33.16_{(0.47)}$ $55.97(\downarrow 1.54)$ $2.286(\uparrow 0.031)$
$137$k $50$ $0.1$ $32$ $50.30_{(0.70)}$ $50.38_{(1.03)}$ $78.80_{(1.29)}$ $66.27_{(1.10)}$ $33.16_{(0.47)}$ $55.78(\downarrow 1.80)$ $2.285(\uparrow 0.030)$
$208$k $75$ $0.1$ $32$ $50.34_{(0.70)}$ $49.20_{(1.03)}$ $80.10_{(1.26)}$ $66.97_{(1.10)}$ $33.19_{(0.47)}$ $55.96(\downarrow 1.62)$ $2.286(\uparrow 0.031)$
@@
$0.1$ 00 $32$ $50.77_{(0.70)}$ $48.95_{(1.03)}$ $80.80_{(1.25)}$ $66.43_{(1.10)}$ $33.16_{(0.47)}$ $56.02(\downarrow 1.56)$ $2.286(\uparrow 0.031)$
$0.15$ 00 $32$ $49.12_{(0.70)}$ $49.66_{(1.03)}$ $81.80_{(1.22)}$ $66.38_{(1.10)}$ $32.84_{(0.47)}$ $55.96(\downarrow 1.62)$ $2.292(\uparrow 0.037)$
$0.2$ 00 $20$ $49.37_{(0.70)}$ $49.87_{(1.03)}$ $79.70_{(1.27)}$ $65.40_{(1.11)}$ $33.08_{(0.47)}$ $55.48(\downarrow 2.10)$ $2.306(\uparrow 0.051)$
@@
Appendix C Experimental Details
C.1 General Setup

### v2

@@ Appendix C Related Works
LLMs are typically trained on a vast amount of data that are rich in knowledge, and extensive studies have investigated how much knowledge LLMs can acquire from the training data. Pioneering studies [petroni2019languagebase, roberts2020muchpack, da2021analyzing] demonstrate that LLMs can capture a substantial amount of knowledge, suggesting their potential as knowledge bases. To quantify the relationship between model size and knowledge storage, allen2024physics and lu2024scaling discover a linear relationship between models’ knowledge capacity and their parameter count by training LLMs on data only containing fixed-format knowledge for sufficiently long horizons. Later, nichani2025understanding formally proved this linear relationship. In contrast, this paper examines the data mixing scenario and demonstrates that this linear scaling can be disrupted when the knowledge-dense dataset is mixed with vast amounts of web-scraped data. Another important factor is the frequency of occurrence for knowledge.
@@
This paper identifies phase transitions in knowledge acquisition within data mixtures with respect to model size and mixing ratio. Some relevant observations can be found in previous papers, but we takes a more direct and systematic approach. kandpal2023large, mallen-etal-2023-trust, sun2024head find that LLMs can perform poorly on low-frequency knowledge. pmlr-v235-ghosal24a-understanding show that frequency of knowledge in the pre-training data determines how well the model encodes the knowledge, which influences its extractability after QA fine-tuning. Taking a more microscopic view, chang2024how insert a few pieces of new knowledge during training and track their loss. By fitting a forgetting curve, they conjecture that the model may fail to learn the knowledge if its frequency is lower than some threshold.
@@
Our findings also relate to prior observations on the memorization and forgetting behaviors of LLMs, but we explicitly characterize phase transitions in the context of data mixing. carlini2023quantifying show that memorization of training data follows a log-linear relationship with model size, the number of repetitions, and prompt length. biderman2024emergent take a data point-level perspective and demonstrate that it is difficult to predict whether a given data point will be memorized using a smaller or partially trained model. By injecting a few new sequences into the training data, huang2024demystifying find that a sequence must be repeated a non-trivial number of times to be memorized. By examining training dynamics, tirumala2022memorizationwooverfitting observe that memorization can occur before overfitting and that larger models memorize faster while forgetting more slowly. zucchet2025language study the training dynamics governing factual knowledge acquisition of LLMs and find that the performance can undergo a plateau before the model acquires precise knowledge, during which the attention-based circuits form. From a theoretical perspective, feldman2020tale prove that memorization of training labels is necessary to achieve near-optimal generalization error for long-tailed data distributions.
@@
LLM performance is significantly influenced by the mixing proportions of the training data from different domains. Our paper is related to a line of studies that optimize the mixing proportions by modeling LLM performance as a function of the mixing proportions [liu2024regmix, kang2024autoscale, ye2024datamixinglaws, ge2024bimix]. However, their datasets can be highly heterogeneous even within a single domain (e.g., OpenWebText, Pile-CC) while we focus on mixing a uniform, knowledge-dense dataset into web-scraped data.
Appendix D Additional Experimental Results
D.1 Additional Results for Phase Transitions
@@
(a) Validation loss on SynBio-1.28M for 410M models.
(b) Validation loss on SynBio-320k for $r=0.1$.
Figure 10: In addition to discrete metrics like accuracy, we can also observe phase transitions in validation loss, a continuous metric.
Table 1: We replicate the experiments in Figure 2(a) with three different random seeds and report the mean and standard deviation below. While accuracy varies slightly with random seeds, the phase transition behavior remains consistent and clearly observable across runs.
$r$ Mean Acc. (%) Std. Dev. (%)
0.1 0.4 0.0
0.15 0.4 0.0
0.2 0.4 0.0
0.25 0.8 0.0
0.3 7.3 2.8
0.35 27.5 2.8
0.4 40.4 3.1
0.45 58.1 3.1
D.2 Ablation Studies
(a) Vary the batch size.
(b) Vary the peak learning rate.
Consistent Trends Across Different Batch Sizes.
Consistent trends across learning rate values and schedules.
D.3 Additional Discussions and Experiments on Reasoning Tasks
(a) Accuracy on Max-over-N for 70M models.
(b) Accuracy on Max-over-N for $r=0.2$.
Figure 12: Phase transitions in the Max-over-N task. Here, we set $N=30$.
Discussion on the slope calculation task: model is indeed learning the procedure rather than memorizing the training data.
We show that the model cannot rely purely on memorization to solve the slope calculation tasks specified in Section 3.3. Specifically, the total number of possible slope calculation problems in our setup is $100\times 100\times 99\times 100=9.9\times 10^{7}$. In contrast, a typical training run in Figure 5 with $r=0.4$ sees fewer than $3.5\times 10^{6}$ unique slope calculation examples, less than 5% of the full space. Despite this limited coverage, the model still achieves 60% accuracy at test time, where examples are uniformly sampled from the full distribution. This substantial generalization beyond the training data suggests that the model is indeed learning a generalizable computation procedure, rather than memorizing specific input-output pairs.
Experiments on Another Reasoning Task with Larger Input Space: Max-over-N.
The slope calculation task suggests that the model learns a generalizable procedure. To test this hypothesis in a more challenging setting, we introduce another task named “Max-over-N”. This task is explicitly designed with a vastly larger input space, rendering memorization computationally infeasible. Specifically, the model is asked to find the maximum number given a list of $N=30$ integers, each randomly sampled from $\{0,1,\cdots,99\}$. This creates an enormous input space of $10^{60}$. The format of the training samples is shown in LABEL:tab:max-over-n-question-template.
Following the setup in Section 3.3, we add 3M tokens of such examples to the OpenWebMath dataset (accounting for less than 0.02% of the original OpenWebMath token count) to create a modified version. We then train Pythia models on the mixture of FineWeb-Edu and the modified OpenWebMath dataset, with $r$ denoting the mixing ratio of the modified OpenWebMath. For evaluation, we generate 1,000 test samples of “Max-over-N” and assess whether the model outputs the correct final answer.
As shown in Figure 12, we observe the same phase transition phenomena with respect to both model size and mixing ratio, consistent with our main findings. The key result is that the 70M model achieves 60% test accuracy after training on fewer than 3,500 unique examples (at $r=0.5$). This training set is an infinitesimal fraction of the $10^{60}$ possible inputs. This clearly demonstrates generalization beyond memorization.
D.4 Additional Results for Validating the Power-Law Relationship of Threshold Frequency and Model Size
In Figure 13, we relax the constraint of training on the same data mixture and investigate the overall trend between model size and $P_{\mathrm{thres}}$. We add the Llama-3 [dubey2024llama3] family, and evaluate both base and instruction-tuned models for all families, totaling 30 models. Interestingly, in Figure 13, $\log$ model size and $\log P_{\mathrm{thres}}$ also exhibit a linear relationship, with most models falling within the 95% confidence interval. We further use models from the OLMo [groeneveld-etal-2024-olmo] family as a validation set, where predictions of the fitted power law closely match the ground truth.
Potential Application: Inferring the Size of Proprietary Models.
D.5 Additional Plots for Strategies to Enhance Knowledge Acquisition
@@
Figure 14: Additional plots for strategies to enhance knowledge acquisition.
D.6 Detailed Performance on SynBio
@@
NN $\rho$ (%) $r$ Birth date Birth city University Major Employer Avg.
@@
NN $\rho$ (%) $r$ Training tokens (B) Birth date Birth city University Major Employer Avg.
@@
D.7 Detailed Downstream Performance
We employ the lm-eval-harness [eval-harness] codebase to evaluate the zero-shot performance on five downstream tasks, including LAMBADA [paperno2016lambada], ARC-E [clark2018arc], PIQA [bisk2020piqa], SciQ [welbl2017sciq], and HellaSwag [zellers2019hellaswag], covering core capabilities such as text understanding, commonsense reasoning, and question answering. We compute the validation loss on about 50M tokens on a holdout set from the Pile or FineWeb-Edu. The detailed downstream performance and validation loss for applying the random subsampling strategy to SynBio and WikiBio are presented in Tables 3 and 4, respectively. Additionally, we report the detailed downstream results for applying CKM to WikiBio in Table 5.
@@ D.7 Detailed Downstream Performance
NN $\rho$ (%) $r$ LAMBADA ARC-E Sciq PIQA HellaSwag Avg. FineWeb-Edu val. loss
- - 0 $38.25_{(0.68)}$ $61.83_{(1.00)}$ $83.60_{(1.17)}$ $68.01_{(1.09)}$ $35.04_{(0.48)}$ $57.35$ $2.667$
1.281.28M $100$ $0.1$ $34.56_{(0.66)}$ $62.33_{(0.99)}$ $83.50_{(1.17)}$ $68.34_{(1.09)}$ $35.13_{(0.48)}$ $56.77(\downarrow 0.58)$ $2.668(\uparrow 0.001)$
1.281.28M $100$ $0.2$ $34.43_{(0.67)}$ $62.13_{(0.99)}$ $83.80_{(1.17)}$ $68.12_{(1.09)}$ $35.39_{(0.48)}$ $56.77(\downarrow 0.58)$ $2.668(\uparrow 0.001)$
1.281.28M $100$ $0.3$ $33.94_{(0.66)}$ $60.77_{(1.00)}$ $80.80_{(1.25)}$ $66.54_{(1.10)}$ $34.23_{(0.47)}$ $55.26(\downarrow 2.09)$ $2.722(\uparrow 0.054)$
320320k $25$ $0.2$ $36.70_{(0.67)}$ $60.35_{(1.00)}$ $(82.70_{1.20})$ $67.74_{(1.09)}$ $34.76_{(0.48)}$ $56.45(\downarrow 0.90)$ $2.686(\uparrow 0.019)$
640640k $50$ $0.2$ $36.58_{(0.67)}$ $60.61_{(1.00)}$ $83.30_{(1.18)}$ $66.65_{(1.10)}$ $34.53_{(0.47)}$ $56.33(\downarrow 1.02)$ $2.688(\uparrow 0.021)$
720720k $56.25$ $0.2$ $35.61_{(0.67)}$ $60.94_{(1.00)}$ $83.00_{(1.19)}$ $67.14_{(1.10)}$ $34.54_{(0.47)}$ $56.25(\downarrow 1.10)$ $2.687(\uparrow 0.020)$
800800k $62.5$ $0.2$ $35.20_{(0.67)}$ $60.48_{(1.00)}$ $83.40_{(1.20)}$ $66.54_{(1.10)}$ $34.45_{(0.47)}$ $56.01(\downarrow 1.34)$ $2.688(\uparrow 0.021)$
@@
NN $\rho$ (%) $r$ Training tokens (B) LAMBADA ARC-E Sciq PIQA HellaSwag Avg. Pile val. loss
@@
2.562.56M $100$ $0.2$ $64$ $53.68_{(0.69)}$ $51.47_{(1.03)}$ $81.00_{(1.24)}$ $68.77_{(1.08)}$ $35.91_{(0.48)}$ $58.17(\downarrow 2.01)$ $2.184(\uparrow 0.016)$
2.562.56M $100$ $0.4$ $24$ $52.38_{(0.70)}$ $51.47_{(1.03)}$ $80.70_{(1.25)}$ $68.17_{(1.09)}$ $34.95_{(0.48)}$ $57.53(\downarrow 2.65)$ $2.198(\uparrow 0.030)$
1.281.28M $50$ $0.2$ $64$ $54.71_{(0.69)}$ $52.86_{(1.02)}$ $81.30_{(1.23)}$ $68.99_{(1.08)}$ $35.48_{(0.48)}$ $58.67(\downarrow 1.51)$ $2.189(\uparrow 0.022)$
@@
NN $\rho$ (%) $r$ Training tokens (B) LAMBADA ARC-E Sciq PIQA HellaSwag Avg. Pile val. loss
@@
277277k $100$ $0.1$ $32$ $50.77_{(0.70)}$ $48.95_{(1.03)}$ $80.80_{(1.25)}$ $66.43_{(1.10)}$ $33.16_{(0.47)}$ $56.02(\downarrow 1.56)$ $2.286(\uparrow 0.031)$
277277k $100$ $0.15$ $32$ $49.12_{(0.70)}$ $49.66_{(1.03)}$ $81.80_{(1.22)}$ $66.38_{(1.10)}$ $32.84_{(0.47)}$ $55.96(\downarrow 1.62)$ $2.292(\uparrow 0.037)$
277277k $100$ $0.2$ 2020 $49.37_{(0.70)}$ $49.87_{(1.03)}$ $79.70_{(1.27)}$ $65.40_{(1.11)}$ $33.08_{(0.47)}$ $55.48(\downarrow 2.10)$ $2.306(\uparrow 0.051)$
6969k $25$ $0.1$ $32$ $48.63_{(0.70)}$ $50.59_{(1.03)}$ $81.00_{(1.24)}$ $66.49_{(1.10)}$ $33.16_{(0.47)}$ $55.97(\downarrow 1.54)$ $2.286(\uparrow 0.031)$
137137k $50$ $0.1$ $32$ $50.30_{(0.70)}$ $50.38_{(1.03)}$ $78.80_{(1.29)}$ $66.27_{(1.10)}$ $33.16_{(0.47)}$ $55.78(\downarrow 1.80)$ $2.285(\uparrow 0.030)$
208208k $75$ $0.1$ $32$ $50.34_{(0.70)}$ $49.20_{(1.03)}$ $80.10_{(1.26)}$ $66.97_{(1.10)}$ $33.19_{(0.47)}$ $55.96(\downarrow 1.62)$ $2.286(\uparrow 0.031)$
@@
$0.1$ 0 $32$ $50.77_{(0.70)}$ $48.95_{(1.03)}$ $80.80_{(1.25)}$ $66.43_{(1.10)}$ $33.16_{(0.47)}$ $56.02(\downarrow 1.56)$ $2.286(\uparrow 0.031)$
$0.15$ 0 $32$ $49.12_{(0.70)}$ $49.66_{(1.03)}$ $81.80_{(1.22)}$ $66.38_{(1.10)}$ $32.84_{(0.47)}$ $55.96(\downarrow 1.62)$ $2.292(\uparrow 0.037)$
$0.2$ 0 2020 $49.37_{(0.70)}$ $49.87_{(1.03)}$ $79.70_{(1.27)}$ $65.40_{(1.11)}$ $33.08_{(0.47)}$ $55.48(\downarrow 2.10)$ $2.306(\uparrow 0.051)$
@@
Appendix E Experimental Details
E.1 General Setup